OpenStudy (anonymous):

Solve negative five times square root of x minus 9 end root minus 7 equals 23

OpenStudy (anonymous):

\[-5 \times \sqrt{x-9} -7 = 23\] like that?

OpenStudy (anonymous):

if so, do no work the left hand side is negative,and the right hand side is positive, so there is no solution

OpenStudy (anonymous):

@gerryliyana yes!!!!

OpenStudy (anonymous):

the possible answers are x=3 x=45 x=-6 none of the above @satellite73

OpenStudy (anonymous):

\[-5 \times \sqrt {x-9} -7 =23\] \[-5 \times \sqrt {x-9} =23 +7\] \[-5 \times \sqrt {x-9} =30\] \[\sqrt{x-9} = \frac{ 30 }{ -5 } = -6\] \[x-9 = (-6)^2 = 36\] \[x= 36 + 9\] \[x = 45\]